ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 251, 2022

Aaronson, J. and T. Sera Tied-down occupation times of infinite ergodic transformations

1

Abért, M. and S. Mellick Point processes, cost, and the growth of rank in locally compact groups

47

Benjamini, I., M. Fraczyk and G. Kun Expander spanning subgraphs with large girth

155

Bergelson, V. and A. Ferré Moragues Juxtaposing combinatorial and ergodic properties of large sets of integers

173

Bowen, L. and P. Burton Locally compact sofic groups  239

Burton, P. See L. Bowen Cyr, V., A. Johnson, B. Kra and A. Şahi^{}İn The complexity

threshold for the emergence of Kakutani inequivalence

271

Downarowicz, T. and G. Zhang Tail variational principle and asymptotic h-expansiveness for amenable group actions

301

Ferré Moragues, A. See V. Bergelson

Foreman, M. and B. Weiss Odometer Based Systems  327

Fraczyk, M. See I. Benjamini Fricker, E. and A. Furman Quasi-Fuchsian vs negative curvature

metrics on surface groups

365

Frisch, J. R., A. S. Kechris and F. Shinko Lifts of Borel actions on quotient spaces

379

Furman, A. See E. Fricker Furstenberg, H. IP -systems and Recurrence in Ergodic Theory: an

Update

423

Ghosh, A., A. Gorodnik and A. Nevo Counting intrinsic Diophantine approximations in simple algebraic groups

443

Glasner, E. On a question of Kazhdan and Yom Din  467

Gorodnik, A. See A. Ghosh Gutman, Y. and Z. Lian Maximal pronilfactors and a topological

Wiener–Wintner theorem

495

Hochman, M. A short proof of Host’s equidistribution theorem  527

Huang, W., S. Shao and X. D. Ye Multiply minimal points for the product of iterates

541

Johnson, A. See V. Cyr Kechris, A. S. See J. R. Frisch Keller, G., M. Lemańczyk, C. Richard and D. Sell On the

Garden of Eden theorem for ℬ-free subshifts

567

Kifer, Yu. Strong diffusion approximation in averaging with dynamical systems fast motions

595

Kra, B. See V. Cyr Kun, G. See I. Benjamini Lemańczyk, M. See G. Keller Levin, G. and F. Przytycki No hyperbolic sets in J_{∞} for infinitely

renormalizable quadratic polynomials

635

Levit, A. and A. Lubotzky Uncountably many permutation stable groups

657

Lian, Z. See Y. Gutman Lubotzky, A. See A. Levit Mellick, S. See M. Abért Nevo, A. See A. Ghosh Przytycki, F. See G. Levin Richard, C. See G. Keller Rühr, R. and O. Sarig Effective intrinsic ergodicity for countable

state Markov shifts

679

Şah^{}İn, A. See V. Cyr Sarig, O. See R. Rühr Sell, D. See G. Keller Sera, T. See J. Aaronson Shao, S. See W. Huang Shinko, F. See J. R. Frisch Tsukamoto, M., M. Tsutaya and M. Yoshinaga G-index, topo-

logical dynamics and the marker property

737

Tsutaya, M. See M. Tsukamoto Weiss, B. See M. Foreman Ye, X. D. See W. Huang Yoshinaga, M. See M. Tsukamoto Zhang, G. See T. Downarowicz